OpenStudy (anonymous):

A right triangle has a hypotenuse of length 15 and a leg of length 9. What is the length of the other leg of the triangle?

OpenStudy (anonymous):

The other leg is 12 units long, am I right?

OpenStudy (anonymous):

@Shane_B

OpenStudy (anonymous):

A^2+b^2 = c^2 first see the variable you know. We know the Hypotenuse (C) and one leg (A or B, we'll say it's A for this example) Now take this and put it into the theorem above and you get 11^2 + B^2 = 15^2 now isolate B which means subtract the square of 11 (121) from the side that B is on and get b^2 = 15^2 + 11^2 which means B^2 = 225 +121 now add like terms (225 and 121) to get B^2 = 346 square root each side to get B = 18.60 (approximately) now you have the length of the other side. Now put what you know into the formula for the area of a triangle which is A = 1/2BH Area equals half the product of B and H. This formula only goes for A right triangle... I think... do the rest yourself good luck. EDIT =------- Sorry I did some of the work wrong but you get the idea, instead of subtracting 121 from 225 I added it but whatever I gave you the formula and operations you should be able to do it yourself :)

OpenStudy (anonymous):

A=83.7?

OpenStudy (anonymous):

@MajicMuzyk

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank you so much for the explanation!

OpenStudy (anonymous):

Your Welcome